There are 5 ways from city A to city B and 4 ways from city B to city C. Find the number n of ways in which a person can travel:

1. From A to C passing through B.

2. From A to C passing through B and then going back.

3. From A to C passing through B. and then going back, using a given way only once.
Answer #1

1)
From A to C passing through B=
  

2)
From A to C passing through B and then going back=20*20=400

3)
From A to C passing through B. and then going back, using a given way only once=5*4*3*4=240